OpenStudy (benlindquist):

3/-x = 7y/x. free medal

OpenStudy (anonymous):

There is no solution to this problem

OpenStudy (ghostgate):

@TypnoEchosis take a look at this: There is a solution to the problem: 3/-x = 7y/x - I believe we would multiply 'x' to both sides: x(3/-x) = 7y - With this, you would multiply x by 3/-x to get: x(3/-x) -> x/1 * 3/-x = 3x/-x -> -3x/x = -3 - Now we have: -3 = 7y - So... with the finalization: -3/7 = y - Which can be worded as: y = -3/7 Now we solve for x: 3/-x = 7(-3/7)/x - Now we multiply 7 by -3/7ths. 3/-x = -3/x - Now since we come to the conclusion that this is true, 'x' has to be 1: x = 1 - The reason being, it's the only thing that makes sense: 3/-1 = -3 and -3/1 = -3. So the answer I believe would be: x = 1 y = -3/7 I'm hoping I did this right, if this was an actual problem I hope it helped in someway! Regardless, have a great day! ;) {---Ghostgate---}

OpenStudy (anonymous):

@Ghostgate That seem's explainable. Thanks for correcting me. :)

OpenStudy (ghostgate):

@TypnoEchosis No problem, I just wanted to point that out. ;)